OpenStudy (anonymous):

Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

myininaya (myininaya):

write cos(2x) in terms of cos^2(x) using an identity then you would have a quadratic type equation

myininaya (myininaya):

recall the double angle identity

OpenStudy (anonymous):

cos^2x + 2 cos x + 1 = 0

myininaya (myininaya):

cos(2x) doesn't equal cos^2(x)

myininaya (myininaya):

use an identity for cos(2x)

OpenStudy (anonymous):

no that's the original problem. the squared didn't show at first.

OpenStudy (anonymous):

I took out a gcf of cosx and was left with cosx=0 and cosx=-3

myininaya (myininaya):

do you know how to solve u^2+2u+1=0?

myininaya (myininaya):

the gcf isn't cos(x) because cos(x) doesn't divide 1

OpenStudy (anonymous):

no I don't , i got 3pi/2 ? but im not sure if its correct

myininaya (myininaya):

do you know how to solve u^2+2u+1=0?

OpenStudy (anonymous):

no

myininaya (myininaya):

so you were never thought how to use the quadratic formula or how to factor?

myininaya (myininaya):

taught*

OpenStudy (anonymous):

yes i have been taught that. its just that we haven't done it in awhile so ts a little blurry

myininaya (myininaya):

that is the way you are expected to solve this one

myininaya (myininaya):

the coefficient of u^2 is 1 so all you need to answer is what two numbers multiply to be c and add up to be b in other words what two numbers multiply to be 1 and add up to be 2?

OpenStudy (anonymous):

1 and 1

myininaya (myininaya):

so u^2+2u+1 can be written as (u+1)(u+1) now back to solving u^2+2u+1=0 so we have (u+1)(u+1)=0 which means u has to be equal to?

OpenStudy (anonymous):

-1

myininaya (myininaya):

so if u=-1 and we solved u^2+2u+1=0 instead of cos^2(x)+2cos(x)+1=0 then that means u=cos(x) which means cos(x)=-1

myininaya (myininaya):

can you solve this equation for x cos(x)=-1?

OpenStudy (anonymous):

ohhh so it would be (pi)

OpenStudy (anonymous):

i understand it now. thank you

myininaya (myininaya):

pi is a solution between 0 and 2 pi there are many more solutions outside that domain of numbers